# IBD-PSC: Input-level Backdoor Detection via Parameter-oriented Scaling Consistency

### v2

@@ 1 Introduction
Backdoor attack is an emerging training-phase threat to deep neural networks (DNNs) (Li et al., 2022). A backdoored model behaves normally on benign samples while misclassifying malicious samples containing the adversary-specified patterns (i.e., triggers). This attack could happen whenever the training stage is not fully controlled. It poses a significant threat to the lifecycle and supply chain of DNNs.
@@
Fortunately, the answer is yes! In this paper, we reveal that the prediction confidences of poisoned samples have parameter-oriented scaling consistency (PSC). Specifically, we scale up the learned parameters of the batch normalization (BN) layers, which are widely exploited in advanced DNN structures. We demonstrate that the prediction confidences of poisoned samples are significantly more consistent than those of benign ones when the number of amplified BN layers increases. In particular, we show that this intriguing phenomenon is not accidental, where we prove that we can always find a scaling factor for BN parameters to expose latent backdoors for all attacked models (under some classical assumptions in learning theory). The scaled model can misclassify benign samples while maintaining the predictions of poisoned samples, leading to PSC phenomenon.
@@ 2.1 Backdoor Attacks
Poison-only Backdoor Attacks. In these attacks, the adversaries can only manipulate the training dataset. Gu et al. (Gu et al., 2017) proposed the first poison-only attack (i.e., BadNets). BadNets poisoned a few training samples by patching a predefined trigger, e.g., a $3\times 3$ white square, onto the bottom right corner of these samples. It then altered the labels of the modified samples to an adversaries-specified target label. Models trained on such poisoned training sets create a relation between the trigger and the target label. Subsequent studies further developed more stealthy attack methods, including invisible and clean-label attacks. The former methods (Chen et al., 2017; Li et al., 2021c) typically used imperceptible triggers to bypass manual detection, while the latter ones  (Turner et al., 2019; Zeng et al., 2023) maintained the ground-truth label of poisoned samples. Besides, there are also the physical attack (Wenger et al., 2021; Gong et al., 2023; Xu et al., 2023) that adopt physical objects or spatial transformations as triggers and the adaptive attack methods (Tang et al., 2021; Qi et al., 2023) that are specifically designed to evade defenses.
Training-controlled Backdoor Attacks. In these attacks, adversaries can modify both the training dataset and the training process. One line of work aimed to circumvent existing defenses and human detection. For instance, the adversaries may introduce a ‘noise mode’ (Nguyen & Tran, 2020, 2021) or incorporate well-designed regularization terms into training loss (Li et al., 2020; Doan et al., 2021; Xia et al., 2022). Another line of work focused on augmenting the effectiveness of attacks. For instance, Wang et al. (Wang et al., 2022b) exploited learning algorithms beyond supervised learning to ensure the correct injections of subtle triggers. Besides, Li et al. (Li et al., 2021c) and Zhang et al. (Zhang et al., 2022) introduced spatial transformations to poisoned samples to hide the triggers more robustly, extending the threat of backdoor attacks to the real physical scenarios.
@@ 2.2 Backdoor Defenses
Based on the stage of the model lifecycle where defense occurs, existing defenses can be mainly divided into five main categories: (1) data purification (Tran et al., 2018; Li et al., 2021b; Jebreel et al., 2023), (2) poison suppression (Wang et al., 2022a; Huang et al., 2022; Tang et al., 2023), (3) model-level backdoor detection (Wang et al., 2019; Xiang et al., 2023; Wang et al., 2024), (4) model-level backdoor mitigation (Liu et al., 2018; Zeng et al., 2022; Guo et al., 2023a), and (5) input-level backdoor detection (IBD) (Gao et al., 2021; Liu et al., 2023; Guo et al., 2023b). Specifically, data purification intends to filter out all poisoned samples in a given (third-party) dataset. It usually needs to train a model before identifying the influence of each training sample; Poison suppression aims to hinder the model’s learning of the poisoned samples by modifying its training process to prevent backdoor creation; Model-level detection usually trains a meta-classifier or approximates trigger generation to determine whether a suspicious model contains hidden backdoors; IBD detects and prevents malicious inputs and acts as a ‘firewall’ of deployed models. In general, the first four strategies demand substantial computational resources since they typically necessitate model training or fine-tuning. However, these resources are unavailable for many researchers and developers, especially those using third-party models. This paper primarily focuses on IBD, which is more computation-friendly.
Previous IBD methods (Chou et al., 2020; Gao et al., 2021; Liu et al., 2023) are effective under certain (implicit) assumptions concerning the backdoor triggers. For example, STRIP (Gao et al., 2021) posited that trigger features play a dominant role, and the predictions of poisoned samples will not be affected even when benign features are overlaid. These assumptions can be easily circumvented by adaptive backdoor attacks (Nguyen & Tran, 2020; Li et al., 2021a). To the best of our knowledge, the most advanced IBD method is SCALE-UP (Guo et al., 2023b). It amplified all pixel values of an input sample with varying intensities and treated it as poisoned if the predictions were consistent. However, SCALE-UP inherited some potential limitations due to pixel value constraints (bounded in [0, 255]). For example, these constraints may alter predictions of poisoned samples, as amplification can transform higher pixel values into the maximum value of 255, causing triggers (e.g., a white square) to disappear. How to design effective yet efficient IBD methods is still a critical open question.
@@
Figure 2: The average confidence of benign and poisoned samples when amplifying the parameters of different number of BN layers under benign and backdoored models (starting from the last layer).
@@ 3 Parameter-oriented Scaling Consistency
Batch Normalization. Let $\phi(\cdot;\gamma,{\bm{\beta}})$ denotes the BN function, for a given batch feature maps ${\bm{a}}$, the BN operation transforms it into $\bm{b}$: $\bm{b}=\phi({\bm{a}};\gamma,{\bm{\beta}})$. This transformation is expressed as $\phi({\bm{a}};\gamma,{\bm{\beta}})=\gamma\left(\frac{{\bm{a}}-{\bm{\mu}}_{a}}{\sqrt{\sigma_{a}^{2}+\epsilon}}\right)+{\bm{\beta}}$, where $\epsilon$ is a small value, ${\bm{\mu}}_{a}$ and $\sigma_{a}$ are the mean and standard deviation of ${\bm{a}}$, respectively. The $\gamma$ and ${\bm{\beta}}$ are learnable parameters, designed to scale and shift normalized features, and learned during the training process.
[…]
Definition 1.
We obtain a parameter-amplified model by performing scalar multiplication on the learned parameters of BN layers with a fixed magnification factor (e.g., 1.5). The scaling process starts from the final BN layer and progressively moves backward to the preceding layers. The PSC is defined as the average prediction confidence (probability) on the originally predicted label over a series of parameter-amplified variants of the deployed model. In particular, we select the label predicted upon the unmodified model as the originally predicted label.
Settings. In this section, we adopt BadNets (Gu et al., 2017), WaNet (Nguyen & Tran, 2021), and BATT (Xu et al., 2023) on CIFAR-10 (Krizhevsky et al., 2009) as examples for analyses. They are the representative of (1) patch-based attack, (2) sample-specific attack, and (3) physical attack, respectively. We exploit a standard ResNet-18 (He et al., 2016a) as our model structure. It contains twenty BN layers. For all attacks, we set the poisoning rate as 10%. Specifically, for each benign and poisoned image, we scale up on the BN parameters (i.e., $\gamma$ and ${\bm{\beta}}$) with $\omega=1.5$ times starting from the last layer and gradually moving forward to more layers. Similar to (Guo et al., 2023b), we also calculate the average confidence defined as the average probability of samples on the label predicted by the original unamplified model. In this paper, confidence refers to the predicted probability assigned to an input sample for a specified label. For instance, if an image of a cat is predicted as the cat label with a probability of 0.9, then the confidence of the input under the cat label is 0.9. More details are in our Appendix B.
@@
Let $F=FC\circ f_{L}\circ\dots\circ f_{1}$ be a backdoored DNN with $L$ hidden layers and FC denotes the fully-connected layers. Let $x$ be an input, $\bm{b}=f_{l}\circ\cdots\circ f_{1}(x)$ be its batch-normalized feature after the $l$-th layer ($1\leq l\leq L$), and $t$ represent the attacker-specified target class. Assume that $\bm{b}$ follows a mixture of Gaussian distributions. Then the following two statements hold: (1) Amplifying the ${\bm{\beta}}$ and $\gamma$ parameters of the $l$-th BN layer can make $\|\tilde{\bm{b}}\|_{2}$ ($\tilde{\bm{b}}$ is the amplified version of $\bm{b}$) arbitrarily large, and (2) There exists a positive constant $M$ that is independent of $\tilde{\bm{b}}$, such that whenever $\|\tilde{\bm{b}}\|_{2}>M$, then $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\tilde{\bm{b}})=t$, even when $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\bm{b})\not=t$
@@
In practice, we find amplifying only a single BN layer may require an unreasonably large amplification factor, and is unstable among different attacks or even BN layers. A detailed exploration of these observations is presented in Appendix C. Fortunately, as shown in Figure 3, amplifying multiple BN layers with a small factor (e.g., 1.5) can also significantly increase the feature norm in the last pre-FC layer and is more stable and robust across different settings. As such, we amplify multiple layers throughout this work.
@@ 4.2 The Overview of IBD-PSC
In general, as shown in Figure 4, our IBD-PSC has two main stages, including (1) model amplification and (2) input detection. In the first stage, we amplify the BN parameters of different layers in the original model to obtain a series of parameter-amplified models. In the second stage, we calculate the PSC value of the suspicious image based on the obtained models and the original one. A larger the PSC value indicates a higher likelihood that the suspicious image is malicious. The technical details are as follows.
@@ 4.3 Model Amplification
Overview. In this stage, we intend to obtain $n$ different parameter-amplified versions of the original model by scaling up the parameters (i.e., $\gamma$ and ${\bm{\beta}}$) of its different BN layers. In particular, we amplify the later parts of the original model. It is motivated by the previous findings that trigger patterns often manifest as complicated features learned by the deeper (convolutional) layers of DNNs, especially for those attacks with elaborate designs (Huang et al., 2022; Jebreel et al., 2023). This finding is also consistent to our observations in Figure 2. Specifically, let $k$ denote the penultimate BN layers in which we scale up in the first parameter-amplified model. For $i$-th amplified model, we scale up the parameters in the last $(k+i-1)$ BN layers with the same scaling factor $\omega$. Let $\mathcal{F}$ denotes the original model, its parameter-amplified version containing $k$ amplified BN layers with scaling factor $\omega$ (i.e., ^$\hat{\mathcal{F}}_{k}^{\omega}$) can be defined as
@@
where $\hat{f}_{i}^{\omega}$ represents the BN layer of the $i$-th hidden layer undergoing an amplification process. It scales the original BN layer’s parameters $\gamma$ and ${\bm{\beta}}$ by a scaling factor ${\omega}$, i.e., $\hat{\gamma}=\omega\cdot\gamma$ and $\hat{{\bm{\beta}}}=\omega\cdot{\bm{\beta}}$. We also conduct ablation studies in Appendices D and E to assess the impact of amplifying BN layers in a forward sequential manner and the impact of amplifying across all BN layers, respectively.
@@ 4.4 Input Detection
where $y^{\prime}=\mathrm{argmax}\left(\mathcal{F}({\bm{x}})\right)$.
After obtaining the PSC value, IBD-PSC assesses whether the input sample is malicious by comparing it to a predefined threshold $T$. If $\textrm{PSC}>T$, it is marked as a poisoned image.
@@ 5.1 Experiment Settings
Attack Baselines. We evaluate the effectiveness of IBD-PSC against thirteen representative backdoor attacks, including 1) BadNets (Gu et al., 2017), 2) Blend (Chen et al., 2017), 3) LC (Turner et al., 2019), 4) ISSBA (Li et al., 2021a), 5) TaCT (Tang et al., 2021), 6) NARCISSUS (Zeng et al., 2023), 7) Adap-Patch (Qi et al., 2023), 8) BATT (Xu et al., 2023), 9) PhysicalBA (Li et al., 2021c), 10) IAD (Nguyen & Tran, 2020), 11) WaNet (Nguyen & Tran, 2021), 12) BPP (Wang et al., 2022b), and 13) SRA (Qi et al., 2022). The first eight attacks are representative of poison-only attacks, while the last one is a model-controlled attack. The remaining four are training-controlled attacks. More details about the attack baselines are in the Appendix G.
Defense Settings. We compare our defense with classical and advanced input-level backdoor defenses, including STRIP (Gao et al., 2021), TeCo (Liu et al., 2023) and SCALE-UP  (Guo et al., 2023b). We implement these defenses using their official codes with default settings. Our IBD-PSC assumes defenders have access to only 100 benign samples, with default settings of $\omega=1.5$, $n=5$, $\epsilon=60\%$, and $T=0.9$. More details can be found in Appendix H.

### v3

@@ 1 Introduction
Backdoor attacks are an emerging training-phase threat to deep neural networks (DNNs) (Li et al., 2022b). A backdoored model behaves normally on benign samples while misclassifying malicious samples containing adversary-specified patterns (i.e., triggers). This attack could happen whenever the training stage is not fully controlled. It poses a significant threat to the lifecycle and supply chain of DNNs.
@@
Fortunately, the answer is yes! In this paper, we reveal that the prediction confidences of poisoned samples have parameter-oriented scaling consistency (PSC). Specifically, we scale up the learned parameters of the batch normalization (BN) layers, which are widely exploited in advanced DNN structures. We demonstrate that the prediction confidences of poisoned samples are significantly more consistent than those of benign ones when the number of amplified BN layers increases. In particular, we show that this intriguing phenomenon is not accidental, where we prove that we can always find a scaling factor for BN parameters to expose latent backdoors for all attacked models (under some classical assumptions in learning theory). The scaled model can misclassify benign samples while maintaining the predictions of poisoned samples, leading to the PSC phenomenon.
@@ 2.1 Backdoor Attacks
Poison-only Backdoor Attacks. In these attacks, the adversaries can only manipulate the training dataset. Gu et al. (Gu et al., 2017) proposed the first poison-only attack (i.e., BadNets). BadNets poisoned a few training samples by patching a predefined trigger, e.g., a $3\times 3$ white square, onto the bottom right corner of these samples. It then altered the labels of the modified samples to an adversaries-specified target label. Models trained on such poisoned training sets create a relation between the trigger and the target label. Subsequent studies further developed more stealthy attack methods, including invisible and clean-label attacks. The former methods (Chen et al., 2017; Li et al., 2021c) typically used imperceptible triggers to bypass manual detection, while the latter ones  (Turner et al., 2019; Zeng et al., 2023; Gao et al., 2023) maintained the ground-truth label of poisoned samples. Besides, there are also the physical attack (Wenger et al., 2021; Gong et al., 2023; Xu et al., 2023) that adopt physical objects or spatial transformations as triggers and the adaptive attack methods (Tang et al., 2021; Qi et al., 2023) that are specifically designed to evade defenses.
Training-controlled Backdoor Attacks. In these attacks, adversaries can modify both the training dataset and the training process. One line of work aimed to circumvent existing defenses and human detection. For instance, the adversaries may introduce a ‘noise mode’ (Nguyen & Tran, 2020, 2021; Mo et al., 2024; Zhang et al., 2024) or incorporate well-designed regularization terms into training loss (Li et al., 2020; Doan et al., 2021; Xia et al., 2022). Another line of work focused on augmenting the effectiveness of attacks. For instance, Wang et al. (Wang et al., 2022c) exploited learning algorithms beyond supervised learning to ensure the correct injections of subtle triggers. Besides, Li et al. (Li et al., 2021c) and Zhang et al. (Zhang et al., 2022) introduced spatial transformations to poisoned samples to hide the triggers more robustly, extending the threat of backdoor attacks to the real physical scenarios.
@@
Recently, a few works exploit backdoor attacks for positive purposes (Li et al., 2022a, c, 2023b; Guo et al., 2023c; Tang et al., 2023; Ya et al., 2024), which are out of our scope.
@@ 2.2 Backdoor Defenses
Based on the stage of the model lifecycle where defense occurs, existing defenses can be mainly divided into five main categories: (1) data purification (Tran et al., 2018; Li et al., 2021b; Jebreel et al., 2023), (2) poison suppression (Wang et al., 2022a; Huang et al., 2022; Tang et al., 2023), (3) model-level backdoor detection (Wang et al., 2019; Xiang et al., 2023; Wang et al., 2024; Yao et al., 2024; Wang et al., 2023, 2022b), (4) model-level backdoor mitigation (Liu et al., 2018; Zeng et al., 2022; Guo et al., 2023a; Li et al., 2024a, b; Xu et al., 2024), and (5) input-level backdoor detection (IBD) (Gao et al., 2021; Liu et al., 2023; Guo et al., 2023b). Specifically, data purification intends to filter out all poisoned samples in a given (third-party) dataset. It usually needs to train a model before identifying the influence of each training sample; Poison suppression aims to hinder the model’s learning of the poisoned samples by modifying its training process to prevent backdoor creation; Model-level detection usually trains a meta-classifier or approximates trigger generation to determine whether a suspicious model contains hidden backdoors; IBD detects and prevents malicious inputs and acts as a ‘firewall’ of deployed models. In general, the first four strategies demand substantial computational resources since they typically necessitate model training or fine-tuning. However, these resources are unavailable for many researchers and developers, especially those using third-party models. This paper primarily focuses on IBD, which is more computation-friendly.
Previous IBD methods (Chou et al., 2020; Gao et al., 2021; Liu et al., 2023) are effective under implicit assumptions concerning the backdoor triggers. For example, STRIP (Gao et al., 2021) posited that trigger features play a dominant role, and the predictions of poisoned samples will not be affected even when benign features are overlaid. These assumptions can be easily circumvented by adaptive backdoor attacks (Nguyen & Tran, 2020; Li et al., 2021a; Duan et al., 2024). To the best of our knowledge, the most advanced IBD method is SCALE-UP (Guo et al., 2023b). It amplified all pixel values of an input sample with varying intensities and treated it as poisoned if the predictions were consistent. However, SCALE-UP inherited some potential limitations due to pixel value constraints (bounded in [0, 255]). For example, these constraints may alter predictions of poisoned samples, as amplification can transform higher pixel values into the maximum value of 255, causing triggers (e.g., a white square) to disappear. How to design effective yet efficient IBD methods is still a critical open question.
@@
Figure 2: The average confidence of benign and poisoned samples when amplifying different numbers of BN layers under benign and backdoored models (starting from the last layer).
@@ 3 Parameter-oriented Scaling Consistency
Batch Normalization. Let $\phi(\cdot;\gamma,{\bm{\beta}})$ denote the BN function, for a given batch feature maps ${\bm{a}}$, the BN operation transforms it into $\bm{b}$: $\bm{b}=\phi({\bm{a}};\gamma,{\bm{\beta}})$. This transformation is expressed as $\phi({\bm{a}};\gamma,{\bm{\beta}})=\gamma\left(\frac{{\bm{a}}-{\bm{\mu}}_{a}}{\sqrt{\sigma_{a}^{2}+\epsilon}}\right)+{\bm{\beta}}$, where $\epsilon$ is a small value, ${\bm{\mu}}_{a}$ and $\sigma_{a}$ are the mean and standard deviation of ${\bm{a}}$, respectively. The $\gamma$ and ${\bm{\beta}}$ are learnable parameters, designed to scale and shift normalized features, and learned during the training process.
Settings. In this section, we adopt BadNets (Gu et al., 2017), WaNet (Nguyen & Tran, 2021), and BATT (Xu et al., 2023) on CIFAR-10 (Krizhevsky et al., 2009) as examples for analyses. They are the representative of (1) patch-based attack, (2) sample-specific attack, and (3) physical attack, respectively. We exploit a standard ResNet-18 (He et al., 2016a) as our model structure. It contains twenty BN layers. For all attacks, we set the poisoning rate as 10%. Specifically, for each benign and poisoned image, we scale up on the BN parameters (i.e., $\gamma$ and ${\bm{\beta}}$) with $\omega=1.5$ times starting from the last layer and gradually moving forward to more layers. Similar to (Guo et al., 2023b), we also calculate the average confidence defined as the average probability of samples on the label predicted by the original unamplified model. More details are in our Appendix B.
@@
Let $F=FC\circ f_{L}\circ\dots\circ f_{1}$ be a backdoored DNN with $L$ hidden layers and FC denotes the fully connected layers. Let $\bm{x}$ be an input, $\bm{b}=f_{l}\circ\cdots\circ f_{1}(\bm{x})$ be its batch-normalized feature after the $l$-th layer ($1\leq l\leq L$), and $t$ represent the attacker-specified target class. Assume that $\bm{b}$ follows a mixture of Gaussian distributions. Then the following two statements hold: (1) Amplifying the ${\bm{\beta}}$ and $\gamma$ parameters of the $l$-th BN layer can make $\|\tilde{\bm{b}}\|_{2}$ ($\tilde{\bm{b}}$ is the amplified version of $\bm{b}$) arbitrarily large, and (2) There exists a positive constant $M$ that is independent of $\tilde{\bm{b}}$, such that whenever $\|\tilde{\bm{b}}\|_{2}>M$, then $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\tilde{\bm{b}})=t$, even when $\arg\max FC\circ f_{L}\circ\dots\circ f_{l+1}(\bm{b})\not=t$.
@@
In practice, we find amplifying only a single BN layer may require an unreasonably large amplification factor and is unstable among different attacks or even BN layers, as demonstrated in Appendix C. Fortunately, as shown in Figure 3, amplifying multiple BN layers with a small factor (e.g., 1.5) can also significantly increase the feature norm in the last pre-FC layer and is more stable across different settings. As such, we amplify multiple layers throughout this work.
@@ 4.2 The Overview of IBD-PSC
In general, as shown in Figure 4, our IBD-PSC has two main stages, including (1) model amplification and (2) input detection. In the first stage, we amplify the BN parameters of different layers in the original model to obtain a series of parameter-amplified models. In the second stage, we calculate the PSC value of the suspicious image based on the obtained models and the original one. A larger PSC value indicates a higher likelihood that the suspicious image is malicious. The technical details are as follows.
@@ 4.3 Model Amplification
Overview. In this stage, we intend to obtain $n$ different parameter-amplified versions of the original model by amplifying the parameters (i.e., $\gamma$ and ${\bm{\beta}}$) of its different BN layers. In particular, we amplify the later parts of the original model. It is motivated by the previous findings that trigger patterns often manifest as complicated features learned by the deeper (convolutional) layers of DNNs, especially for those attacks with elaborate designs (Huang et al., 2022; Jebreel et al., 2023). This finding is consistent with our observations in Figure 2. Specifically, let $k$ denote the penultimate BN layers in which we scale up in the first parameter-amplified model. For the $i$-th amplified model, we scale up the parameters in the last $(k+i-1)$ BN layers with the same scaling factor $\omega$. Let $\mathcal{F}$ denotes the original model, its parameter-amplified version containing $k$ amplified BN layers with scaling factor $\omega$ (i.e., ^$\hat{\mathcal{F}}_{k}^{\omega}$) can be defined as
@@
where $\hat{f}_{i}^{\omega}$ represents the BN layer of the $i$-th hidden layer undergoing an amplification process. It scales the original BN layer’s parameters $\gamma$ and ${\bm{\beta}}$ by a scaling factor ${\omega}$, i.e., $\hat{\gamma}=\omega\cdot\gamma$ and $\hat{{\bm{\beta}}}=\omega\cdot{\bm{\beta}}$. We also conduct ablation studies in Appendix D and Appendix E to assess the impact of amplifying BN layers in a forward sequential manner and that of amplifying all BN layers, respectively.
@@ 4.4 Input Detection
where $y^{\prime}=\mathrm{argmax}\left(\mathcal{F}({\bm{x}})\right)$. After obtaining the PSC value, IBD-PSC assesses whether the input sample is malicious by comparing it to a predefined threshold $T$. If $\textrm{PSC}>T$, it is marked as a poisoned image.
@@ 5.1 Experiment Settings
Attack Baselines. We evaluate the effectiveness of IBD-PSC against thirteen representative backdoor attacks, including 1) BadNets (Gu et al., 2017), 2) Blend (Chen et al., 2017), 3) LC (Turner et al., 2019), 4) ISSBA (Li et al., 2021a), 5) TaCT (Tang et al., 2021), 6) NARCISSUS (Zeng et al., 2023), 7) Adap-Patch (Qi et al., 2023), 8) BATT (Xu et al., 2023), 9) PhysicalBA (Li et al., 2021c), 10) IAD (Nguyen & Tran, 2020), 11) WaNet (Nguyen & Tran, 2021), 12) BPP (Wang et al., 2022c), and 13) SRA (Qi et al., 2022). The first eight attacks are representative of poison-only attacks, while the last one is a model-controlled attack. The remaining four are training-controlled attacks. More details about the attack baselines are in the Appendix G.
Defense Settings. We compare our defense with classical and advanced input-level backdoor defenses, including STRIP (Gao et al., 2021), TeCo (Liu et al., 2023) and SCALE-UP (Guo et al., 2023b). We implement these defenses using their official codes with default settings. Our IBD-PSC defense maintains a consistent hyper-parameter setting across various attacks and datasets. Specifically, we set $\omega=1.5$, $n=5$, $\epsilon=60\%$, and $T=0.9$. Defenders can only access 100 benign samples as their local samples. More setting details about baseline methods are in Appendix H.